40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 11635     Entry time: Tue Sep 22 16:52:36 2015
Author: ericq 
Type: Summary 
Category: General 
Subject: Random Notes 

Some things bouncing around my head that haven't made it to ELOG yet:

  • Last week, Rana and I were investigating excess power line noise coming from the DFD demodulation. We put transformers on the green beat signals where they arrive at the LSC rack, to avoid connecting their signal ground from the PSL table to the LSC rack ground. This didn't help; it's unclear what the culprit is; maybe the demod board power board?
  • Lately, when the interferometer loses lock, the Y arm will not lock on POY, or even flash its IR resonance. for a little while. The green beam can be locked, the X arm can be locked, and no excess angular noise is evident from glancing at the oplev XY plots. Mysterious.  
  • Sometimes, when writing new values to the C1LSC SDF table, the c1lscepics process dies (though the write is successful). This is highly annoying. This may have been adressed in some slightly newer RCG code. 
  • C1OAF is running with a big red NO SYNC message on its GDS screen. C1LSC has shown this too, but I think only when the SDF/epics crash happens. 
  • C1OAF also doesn't seem to properly load the "safe" SDF table when starting up, and errantly puts ones in every element in the static FF matrix. Be careful when restarting OAF!
ELOG V3.1.3-